# Degeneracy slopes, boundary slopes and exceptional surgery slopes

### v2

@@ Abstract.
This work was partially supported by JSPS KAKENHI Grant Number JP22K03301.
@@ 1. Introduction
Let us start with fundamental terminology used in the paper. See [Rol76] as a basic reference, for example. As usual, by a slope, we mean the isotopy class of a non-trivial unoriented simple closed curve on a torus. The distance between two slopes is defined as the minimal geometric intersection number of their representatives. We use $\Delta(\gamma,\gamma^{\prime})$ to denote the distance between the slopes $\gamma$ and $\gamma^{\prime}$. When a meridian-longitude system on a torus is fixed, slopes on the torus are parametrized by rational numbers (irreducible fractions) with $1/0$ in the usual way, and are identified accordingly. In the following, when the torus is the boundary of the regular neighborhood of a knot in the 3-sphere, or in an integral homology sphere, we always adopt the standard meridian–preferred longitude system. In this case, the distance between two slopes corresponding to $p/q$ and $r/s$ is calculated as $\Delta(p/q,r/s)=|ps-qr|$.
@@
A codimension one foliation $\lambda$ of a closed subset of a 3-manifold $M$ is called a lamination in $M$. The complement of a lamination $\lambda$ decomposes into connected components called complementary regions. A lamination is called essential if it contains no spherical or Reeb torus leaf and if its complement contains neither essential disks, monogons, nor spheres. See [GO89, Section 5] and [Cal07, Chapter 6].
Based on [Bri98] and [Gab97, Definition 8.5], we extend the definition of a degeneracy slope of an essential lamination in a knot exterior to one in a compact orientable 3-manifold $M$ with torus boundary $T$ as follows. See also [Ni20, Subsection 2.1]. Suppose a meridian–longitude system on $\partial M$ is fixed. For an essential lamination $\lambda$ in $M$, (if possible) the degeneracy locus $d(\lambda)$ of $\lambda$ is defined as $n(p,q)$ if there exist $n$ pairwise disjoint, properly embedded, non-parallel annuli in a closed complementary region such that one boundary component lies in a leaf of $\lambda$ and the other is a $p/q$ curve on $\partial M$. Furthermore, $n$ is set to be the maximal number of such annuli. The rational number $p/q$ is called a degeneracy slope of $\lambda$. We remark that this definition is slightly different from that in [Cal07, Chapter 6].
@@
One is [Gab97, Theorem 8.8], stated as follows. Let $K$ be a knot in the 3-sphere $S^{3}$ and $\lambda$ an essential lamination in the exterior $E(K)$. If $d(\lambda)=n(1,0)$, then $1\leq n\leq 4g(K)-2$; otherwise, $d(\lambda)=1(p,1)$ and $|p|\leq 4g(K)-2$, where $g(K)$ denotes the genus of $K$, that is, the genus of the minimal genus Seifert surface of $K$. Note that if $d(\lambda)=1(p,1)$, then $|p/1|=|p|$ is the distance between the degeneracy slope and the longitudinal slope, that is, the boundary slope of a minimal genus Seifert surface of genus $g(K)$, which is known to be essential. Thus, in the case $d(\lambda)=1(p,1)$, the inequality can be rewritten as $\Delta(\delta,\beta)\leq 4g(K)-2$ for a degeneracy slope $\delta=p/1$ and the boundary slope $\beta$ of a minimal genus Seifert surface.
A closed essential surface in a 3-manifold is regarded as an essential lamination with a single compact leaf. In this case, the following result is shown in [IO02]. Let $K$ be a knot in $S^{3}$ with exterior $E(K)$, $F$ an essential surface properly embedded in $E(K)$, and $S$ a closed essential surface in $E(K)$ with an annulus $A$ connecting $S$ and $\partial E(K)$. Let $\beta$ denote the boundary slope of $F$ and $\delta$ the slope on $\partial E(K)$ determined by $A$. Then, $\Delta(\beta,\delta)\leq 2(-\chi(F))/|\partial F|$ holds. Note that if we regard $S$ as an essential lamination with a single compact leaf, the slope $\delta$ is a degeneracy slope of $S$. Also note that if the surface $F$ is orientable and has a single boundary component, then we have $2(-\chi(F))/|\partial F|=4g(F)-2$.
On the other hand, as with train tracks, laminations can be carried by branched surfaces, which allows one to study them combinatorially. In the proof of [Wu98, Theorem 2.5], the following is essentially shown. Suppose $M$ is a hyperbolic manifold with torus boundary $T$. Let $B$ be an essential branched surface in $M$ with a regular neighborhood $N(B)$. (See the next section for details.) Suppose that $M-\mathrm{Int}\,N(B)$ contains an essential annulus $A$ with one boundary on $\partial_{h}N(B)$ and the other a curve on $T$ of slope $\delta$. Then a boundary slope $\beta$ of an essential surface of genus one (punctured torus) satisfies $\Delta(\delta,\beta)\leq 2$. Note that the slope $\delta$ corresponds to a degeneracy slope of an essential lamination carried by $B$.
@@
Here, an essential lamination in a $3$-manifold is called very full if every complementary region is homeomorphic to one of the following two types: (i) an ideal polygon bundle over the circle $S^{1}$, where the polygon has at least two ideal vertices; or (ii) a one-holed ideal polygon bundle over $S^{1}$, where the polygon has at least one ideal vertex. See [Cal07, Definition 6.42] for more details. It is proved by Gabai and Mosher that any compact, orientable, irreducible, atoroidal $3$-manifold with incompressible torus boundary, in particular, a hyperbolic knot complement, contains a very full lamination. See [Mos96, Theorem A] and [Cal07, Theorem 6.47 (Gabai–Mosher)].
@@
It follows from [Men84, Theorem 2 (Meridian Lemma)] and [IO00, Theorem 1] that if $K$ is a prime alternating knot having a closed incompressible surface $S$ in its exterior $E(K)$, then there always exist embedded annuli in $E(K)$ connecting $S$ to $\partial E(K)$, and the slope on $\partial E(K)$ determined by the annuli is meridional. This means that, when we regard $S$ as an essential lamination with a single compact leaf, the degeneracy slope must be meridional.
For 2-bridge knots, which are well known to be alternating knots, the following was shown in [GK90, Theorem 8]. Let $K$ be a fibered 2-bridge knot in $S^{3}$. The stable lamination transverse to the fibration is known to be very full. Its degeneracy locus is then of the form 2(m1)(1,0)2(m-1)\,(1,0), where a positive integer $m$ is determined by the Murasugi sum decomposition of the fiber surface. In particular, the degeneracy slope is meridional.
It was raised as a conjecture in [GK90, Conjecture 13] that if $K$ is a non-torus fibered alternating knot, then $d(\lambda)=n(1,0)$ holds for the stable lamination $\lambda$ transverse to the fibration and some integer $n>1$.
@@
This result was recently proved in [BGH24, Corollary 6.18], by using $\mathrm{Homeo}_{+}(S^{1})$-representations of link groups. This theorem provides an alternative proof. See [BGH24] for details.
@@
For a hyperbolic alternating fibered knot $K$, the stable lamination $\lambda$ transverse to the fibration is known to be very full. Let $d$ be the degeneracy slope for $\lambda$. Then, by [HKM07, Proposition 3.1], the monodromy of the fibration is right-veering (resp. left-veering) if and only if $1/d>0$ (resp. $1/d<0$). Thus, the corollary follows from the theorem above. ∎
This result was recently shown in [BNS25] by using knot Floer homology. The corollary above provides an alternative proof. See [BNS25] for details.
@@
Based on Hatcher-Thurston [HT85] and Hatcher-Oertel [HO89], the following results were shown in [IM07, Theorem 1.4] about boundary slopes for Montesinos knots in $S^{3}$. (See [IM07] for details.) Let $K$ be a Montesinos knot of length at least three. (i) Let $F$ be an essential surface of genus $g\geq 2$ properly embedded in the exterior of $K$ with boundary slope $p/q$. Then, $|q|\leq(-\chi(F))/|\partial F|+1$ holds. In particular, if $F$ is orientable, then $q\leq\max\{g+1,2g-1\}$ holds ([IM07, Theorem 1.1]). (ii) Let $F_{i}$ be essential surfaces of genus $g_{i}\geq 2$ properly embedded in the exterior of $K$ with non-meridional boundary slopes $r_{i}$ for $i=1,2$, respectively. Then, $|r_{1}-r_{2}|\leq 4(g_{1}+g_{2})$ holds ([IM07, Theorem 1.4]).
@@
See the following for related results: [GL87, Theorem 2], [Tor96, Theorem 2, 3], [MT92, Theorem 2], and [MS04, Theorem 1.2] about bounds on denominators of boundary slopes for knots, and also [Rie00, Theorem 5.2], [Tor96, Theorem 1], and [Ich25, Corollary 4.4] about bounds on the distances between boundary slopes for knots.
@@
Finally, we give an application to the study of exceptional surgeries on hyperbolic knots. See, for example, [Boy02] for a survey.
@@
The well-known Hyperbolic Dehn Surgery Theorem due to Thurston [Thu82, Theorem 5.8.2] states that each hyperbolic knot (i.e., a knot with hyperbolic complement) admits only finitely many Dehn surgeries yielding non-hyperbolic manifolds. In view of this, such finitely many exceptions are called exceptional surgeries.
Regarding exceptional surgeries on knots in $S^{3}$, there are two known conjectures. One is about the denominators of exceptional surgery slopes: Every non-trivial exceptional surgery slope $p/q$ for a hyperbolic knot in $S^{3}$ satisfies $|q|\leq 2$. See [Kir97, Problem 1.77 A(3)] for more explanation. Another states: Every non-trivial exceptional surgery slope $r$ for a hyperbolic knot in $S^{3}$ satisfies $|r|\leq 4g(K)$.
The next theorem shows that one of the conjectures always holds.
@@
Similar results were obtained in [Ni20, Theorem 1.1] and [BNS25, Corollary 1.12] for hyperbolic fibered knots in $S^{3}$.
By virtue of the Geometrization of 3-manifolds established in [Per02, Per03b, Per03a], exceptional surgeries are classified into those yielding reducible, toroidal, or (small) Seifert fibered manifolds. Among these, surgeries yielding spherical manifolds, that is, 3-manifolds with finite fundamental groups, in particular lens spaces, have been extensively studied. For these, we have the following.
@@
Let $K$ be a hyperbolic knot in $S^{3}$ and $b$ a non-meridional boundary slope of an orientable essential surface $F$ embedded in the exterior $E(K)$ of $K$. Suppose that the degeneracy locus of a very full essential lamination in $E(K)$ is not $1(1,0)$. Then, if the 3-manifold $K(r)$ obtained by $r$-surgery on $K$ is a lens space or $\pi_{1}(K(r))$ is finite, then $|r-b|\leq 4g(F)-1$, in particular, $|r|\leq 4g(K)-1$.
@@
Let $\gamma=r$ be a slope such that $K(r)$ is a lens space or $\pi_{1}(K(r))$ is finite. We consider the degeneracy slope $\delta$ of a very full essential lamination $\lambda$ in the exterior $E(K)$ of $K$. As stated before, by [Mos96, Theorem A], there always exists a very full essential lamination in $E(K)$, and so the degeneracy slope $\delta$ must exist. See also [Cal07, Theorem 6.47. (Gabai–Mosher)].
When $\delta=1/0$, equivalently, the corresponding degeneracy locus is $n(1,0)$, if $n\geq 2$, the lamination $\lambda$ remains essential in the manifold obtained by surgery along any $\gamma\neq\delta=1/0$ by [GO89]. Moreover, for any $\delta$, if $\Delta(\delta,\gamma)\geq 2$, $\lambda$ remains essential in the manifold obtained by surgery along $\gamma$ by [Cal07, Theorem 6.48. (Gabai–Mosher)].
Since the universal cover of a compact orientable 3-manifold contains an essential lamination is homeomorphic to $\mathbb{R}^{3}$ ([GO89, Theorem 6.1]), in particular, the fundamental group of the manifold must be infinite, any lens space or a 3-manifold with finite fundamental group cannot contain an essential lamination. This implies that, under the assumption that the degeneracy locus is not $1(1,0)$, $\delta\neq 1/0$ and $\Delta(\delta,\gamma)\leq 1$.
@@
If we consider a minimal genus Seifert surface, which is shown to be essential, as $F$, we obtain $|r|\leq 4g(K)-1$. ∎
See the following for related results; [GL87, Theorem 1], [BZ96, Theorem 1.1], [CGLS87], and [GL95, GL00] about the bounds on the denominators of exceptional surgery slopes, and also, [GT00, Theorem 1.1], [KMOS07, Corollary 8.5], [Ras04, Theorem 1] (for les space surgery), [MS03, Proposition 1.3 (i)] (for reducible surgery), [Ter03, Theorem 1.2, Theorem 1.3], [EMnGT19, Theorem 1], [IT03, IT05] (for toroidal surgery), and also [Ich01a, Ich01b], [IO02, Theorem 5.1(B)], about the bounds on the exceptional surgery slopes in terms of the genera of knots.

### v3

@@ 1. Introduction
Let us start with fundamental terminology used in the paper. See [44] as a basic reference, for example. As usual, by a slope, we mean the isotopy class of a non-trivial unoriented simple closed curve on a torus. The distance between two slopes is defined as the minimal geometric intersection number of their representatives. (Note that this is not a distance function on the set of slopes.) We use $\Delta(\gamma,\gamma^{\prime})$ to denote the distance between the slopes $\gamma$ and $\gamma^{\prime}$. When a meridian-longitude system on a torus is fixed, slopes on the torus are parametrized by rational numbers (irreducible fractions) with $1/0$ in the usual way, and are identified accordingly. In the following, when the torus is the boundary of the regular neighborhood of a knot in the 3-sphere, or in an integral homology sphere, we always adopt the standard meridian–preferred longitude system. In this case, the distance between two slopes corresponding to $p/q$ and $r/s$ is calculated as $\Delta(p/q,r/s)=|ps-qr|$.
@@
A codimension one foliation $\lambda$ of a closed subset of a 3-manifold $M$ is called a lamination in $M$. The complement of a lamination $\lambda$ decomposes into connected components called complementary regions. A lamination is called essential if it contains no spherical or Reeb torus leaf and if its complement contains neither essential disks, monogons, nor spheres. See [13, Section 5] and [8, Chapter 6].
Based on [6] and [14, Definition 8.5], we extend the definition of a degeneracy slope of an essential lamination in a knot exterior to one in a compact orientable 3-manifold $M$ with torus boundary $T$ as follows. See also [38, Subsection 2.1]. Suppose a meridian–longitude system on $\partial M$ is fixed. For an essential lamination $\lambda$ in $M$, (if possible) the degeneracy locus $d(\lambda)$ of $\lambda$ is defined as $n(p,q)$ if there exist $n$ pairwise disjoint, properly embedded, non-parallel annuli in a closed complementary region such that one boundary component lies in a leaf of $\lambda$ and the other is a $p/q$ curve on $\partial M$. Furthermore, $n$ is set to be the maximal number of such annuli. The rational number $p/q$ is called a degeneracy slope of $\lambda$. We remark that this definition is slightly different from that in [8, Chapter 6].
@@
One is [14, Theorem 8.8], stated as follows. Let $K$ be a knot in the 3-sphere $S^{3}$ and $\lambda$ an essential lamination in the exterior $E(K)$. If $d(\lambda)=n(1,0)$, then $1\leq n\leq 4g(K)-2$; otherwise, $d(\lambda)=1(p,1)$ and $|p|\leq 4g(K)-2$, where $g(K)$ denotes the genus of $K$, that is, the genus of the minimal genus Seifert surface of $K$. Note that if $d(\lambda)=1(p,1)$, then $\Delta(p/1,0/1)=p$ is the distance between the degeneracy slope and the longitudinal slope, that is, the boundary slope of a minimal genus Seifert surface of genus $g(K)$, which is known to be essential. Thus, in the case $d(\lambda)=1(p,1)$, the inequality can be rewritten as $\Delta(\delta,\beta)\leq 4g(K)-2$ for a degeneracy slope $\delta=p/1$ and the boundary slope $\beta$ of a minimal genus Seifert surface.
A closed essential surface in a 3-manifold is regarded as an essential lamination with a single compact leaf. In this case, the following result is shown in [25]. Let $K$ be a knot in $S^{3}$ with exterior $E(K)$, $F$ an essential surface properly embedded in $E(K)$, and $S$ a closed essential surface in $E(K)$ with an annulus $A$ connecting $S$ and $\partial E(K)$. Let $\beta$ denote the boundary slope of $F$ and $\delta$ the slope on $\partial E(K)$ determined by $A$. Then, $\Delta(\beta,\delta)\leq 2(-\chi(F))/|\partial F|$ holds. Note that if we regard $S$ as an essential lamination with a single compact leaf, the slope $\delta$ is a degeneracy slope of $S$. Also note that if the surface $F$ is orientable and has a single boundary component, then we have $2(-\chi(F))/|\partial F|=4g(F)-2$.
On the other hand, as with train tracks, laminations can be carried by branched surfaces, which allows one to study them combinatorially. In the proof of [48, Theorem 2.5], the following is essentially shown. Suppose $M$ is a hyperbolic manifold with torus boundary $T$. Let $B$ be an essential branched surface in $M$ with a regular neighborhood $N(B)$. (See the next section for details.) Suppose that $M-\mathrm{Int}\,N(B)$ contains an essential annulus $A$ with one boundary on $\partial_{h}N(B)$ and the other a curve on $T$ of slope $\delta$. Then a boundary slope $\beta$ of an essential surface of genus one (punctured torus) satisfies $\Delta(\delta,\beta)\leq 2$. Note that the slope $\delta$ corresponds to a degeneracy slope of an essential lamination if $B$ carries a (necessarily essential) lamination.
@@
Here, an essential lamination in a $3$-manifold is called very full if every complementary region is homeomorphic to one of the following two types: (i) an ideal polygon bundle over the circle $S^{1}$, where the polygon has at least two ideal vertices; or (ii) a one-holed ideal polygon bundle over $S^{1}$, where the polygon has at least one ideal vertex. See [8, Definition 6.42] for more details. It is proved by Gabai and Mosher that any compact, orientable, irreducible, atoroidal $3$-manifold with incompressible torus boundary, in particular, a hyperbolic knot complement, contains a very full lamination. See [37, Theorem A] and [8, Theorem 6.47 (Gabai–Mosher)].
@@
It follows from [35, Theorem 2 (Meridian Lemma)] and [24, Theorem 1] that if $K$ is a prime alternating knot having a closed incompressible surface $S$ in its exterior $E(K)$, then there always exist embedded annuli in $E(K)$ connecting $S$ to $\partial E(K)$, and the slope on $\partial E(K)$ determined by the annuli is meridional. This means that, when we regard $S$ as an essential lamination with a single compact leaf, the degeneracy slope must be meridional.
For 2-bridge knots, which are well known to be alternating knots, the following was shown in [12, Theorem 8]. Let $K$ be a fibered 2-bridge knot in $S^{3}$. The stable lamination transverse to the fibration is known to be very full. Its degeneracy locus is then of the form 2(m1)(1,0)2(m-1)\,(1,0), where the positive integer $m$ is determined by the Murasugi sum decomposition of the fiber surface. In particular, the degeneracy slope is meridional.
It was raised as a conjecture in [12, Conjecture 13] that if $K$ is a non-torus fibered alternating knot, then $d(\lambda)=n(1,0)$ holds for the stable lamination $\lambda$ transverse to the fibration and some integer $n>1$.
@@
The Gabai’s conjecture that degeneracy slopes of fibered alternating knots are meridional was recently proved in [4, Corollary 6.18], by using $\mathrm{Homeo}_{+}(S^{1})$-representations of link groups. This theorem provides an alternative proof. See [4] for details.
@@
For a hyperbolic alternating fibered knot $K$, the stable lamination $\lambda$ transverse to the fibration is known to be very full. Let $d$ be the degeneracy slope for $\lambda$. Then, by [21, Proposition 3.1], the monodromy of the fibration is right-veering (resp. left-veering) if and only if $1/d>0$ (resp. $1/d<0$). Thus, the corollary follows from the theorem above. ∎
This result was recently shown in [2] by using knot Floer homology. The corollary above provides an alternative proof. See [2] for details.
@@
Based on Hatcher-Thurston [20] and Hatcher-Oertel [19], the following results were shown in [22, Theorem 1.4] about boundary slopes for Montesinos knots in $S^{3}$. (See [22] for details.) Let $K$ be a Montesinos knot of length at least three. (i) Let $F$ be an essential surface of genus $g\geq 2$ properly embedded in the exterior of $K$ with boundary slope $p/q$. Then, $|q|\leq(-\chi(F))/|\partial F|+1$ holds. In particular, if $F$ is orientable, then $q\leq\max\{g+1,2g-1\}$ holds ([22, Theorem 1.1]). (ii) Let $F_{i}$ be essential surfaces of genus $g_{i}\geq 2$ properly embedded in the exterior of $K$ with non-meridional boundary slopes $r_{i}$ for $i=1,2$, respectively. Then, $|r_{1}-r_{2}|\leq 4(g_{1}+g_{2})$ holds ([22, Theorem 1.4]).
@@
See the following for related results: [16, Theorem 2], [47, Theorem 2, 3], [36, Theorem 2], and [33, Theorem 1.2] about bounds on denominators of boundary slopes for knots, and also [43, Theorem 5.2], [47, Theorem 1], and [30, Corollary 4.4] about bounds on the distances between boundary slopes for knots.
@@
Finally, we give an application to the study of exceptional surgeries on hyperbolic knots. See, for example, [5] for a survey.
@@
The well-known Hyperbolic Dehn Surgery Theorem due to Thurston [46, Theorem 5.8.2] states that each hyperbolic knot (i.e., a knot with hyperbolic complement) admits only finitely many Dehn surgeries yielding non-hyperbolic manifolds. In view of this, such finitely many exceptions are called exceptional surgeries.
Regarding exceptional surgeries on knots in $S^{3}$, there are two conjectures, both are still open. One is about the denominators of exceptional surgery slopes: Every non-trivial exceptional surgery slope $p/q$ for a hyperbolic knot in $S^{3}$ satisfies $|q|\leq 2$. See [31, Problem 1.77 A(3)] for more explanation. Another states: Every non-trivial exceptional surgery slope $r$ for a hyperbolic knot in $S^{3}$ satisfies $|r|\leq 4g(K)$.
The next theorem shows that one of the conjectures holds for each hyperbolic knot in an integral homology sphere.
@@
Note that it is not shown that at least one of the above statements always holds, i.e., it could be that for some knots the first statement holds, and for some knots the second, but no statement holds for all knots.
Similar results were obtained in [38, Theorem 1.1] and [2, Corollary 1.12] for hyperbolic fibered knots in $S^{3}$.
By virtue of the Geometrization of 3-manifolds established in [39, 41, 40], exceptional surgeries are classified into those yielding reducible, toroidal, or (small) Seifert fibered manifolds. Among these, surgeries yielding spherical manifolds, that is, 3-manifolds with finite fundamental groups, in particular lens spaces, have been extensively studied. For these, we have the following.
@@
Let $K$ be a hyperbolic knot in $S^{3}$ and $b$ a non-meridional boundary slope of an orientable essential surface $F$ embedded in the exterior $E(K)$ of $K$. Suppose that the degeneracy locus of a very full essential lamination in $E(K)$ is not $1(1,0)$. Then, if the 3-manifold $K(r)$ obtained by $r$-surgery on $K$ for a non-meridional slope $r$ is a lens space or $\pi_{1}(K(r))$ is finite, then $|r-b|\leq 4g(F)-1$, in particular, $|r|\leq 4g(K)-1$.
@@
Let $\gamma=r$ be a slope such that $K(r)$ is a lens space or $\pi_{1}(K(r))$ is finite. We consider the degeneracy slope $\delta$ of a very full essential lamination $\lambda$ in the exterior $E(K)$ of $K$. As stated before, by [37, Theorem A], there always exists a very full essential lamination in $E(K)$, and so the degeneracy slope $\delta$ must exist. See also [8, Theorem 6.47. (Gabai–Mosher)].
When $\delta=1/0$, equivalently, the corresponding degeneracy locus is $n(1,0)$, if $n\geq 2$, the lamination $\lambda$ remains essential in the manifold obtained by surgery along any $\gamma\neq\delta=1/0$ by [13]. Moreover, for any $\delta$, if $\Delta(\delta,\gamma)\geq 2$, $\lambda$ remains essential in the manifold obtained by surgery along $\gamma$ by [8, Theorem 6.48. (Gabai–Mosher)].
Since the universal cover of a compact orientable 3-manifold containing an essential lamination is homeomorphic to $\mathbb{R}^{3}$ ([13, Theorem 6.1]), in particular, the fundamental group of the manifold must be infinite, any lens space or a 3-manifold with finite fundamental group cannot contain an essential lamination. This implies that, under the assumption that the degeneracy locus is not $1(1,0)$, $\delta\neq 1/0$ and $\Delta(\delta,\gamma)\leq 1$.
@@
If we consider a minimal genus Seifert surface, which is known to be essential, as $F$, we obtain $|r|\leq 4g(K)-1$. ∎
See the following for related results; [16, Theorem 1], [3, Theorem 1.1], [9], and [17, 18] about the bounds on the denominators of exceptional surgery slopes, and also, [15, Theorem 1.1], [32, Corollary 8.5], [42, Theorem 1] (for les space surgery), [34, Proposition 1.3 (i)] (for reducible surgery), [45, Theorem 1.2, Theorem 1.3], [11, Theorem 1], [26, 27] (for toroidal surgery), and also [28, 29], [25, Theorem 5.1(B)], about the bounds on the exceptional surgery slopes in terms of the genera of knots.